OpenStudy (anonymous):

The sum of two consecutive even integers is –46. What are the numbers?

OpenStudy (anonymous):

do you mean positive 46?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

make a guess and see

OpenStudy (anonymous):

A. –26 and –20 B. –22 and –20 C. –26 and –24 D. –24 and –22 These are the given answers... it is suposed to be -46

OpenStudy (anonymous):

for example, i guess \(-20\) and \(-22\) but they add to \(-42\) so that is not right

OpenStudy (anonymous):

try \(-22\) and \(-24\)

OpenStudy (anonymous):

Lets see..

OpenStudy (anonymous):

@satellite73 THANKS SO MUCH THAT WAS CORRECT!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

x = one number x + 2 = the consecutive even number x+x+2 = -46 2x+2 = -46 2x = -48 x=-24 the first number is -24 the second is -24 +2 =-22 the answer is d) -24, -22

OpenStudy (anonymous):

thanks for the help guys!!